# These authors contributed equally to this work. [1]Pasquale Claudio Africa

### v1

@@ Abstract
The complexity of the cardiovascular system needs to be accurately reproduced in order to promptly acknowledge health conditions; to this aim, advanced multifidelity and multiphysics numerical models are crucial. On one side, Full Order Models (FOMs) deliver accurate hemodynamic assessments, but their high computational demands hinder their real-time clinical application. In contrast, ROMs provide more efficient yet accurate solutions, essential for personalized healthcare and timely clinical decision-making. In this work, we explore the application of computational fluid dynamics (CFD) in cardiovascular medicine by integrating FOMs with ROMs for predicting the risk of aortic aneurysm growth and rupture. Wall Shear Stress (WSS) and the Oscillatory Shear Index (OSI), sampled at different growth stages of the abdominal aortic aneurysm, are predicted by means of Graph Neural Networks (GNNs). GNNs exploit the natural graph structure of the mesh obtained by the Finite Volume (FV) discretization, taking into account the spatial local information, regardless of the dimension of the input graph. Our experimental validation framework yields promising results, confirming our method as a valid alternative that overcomes the curse of dimensionality.
@@ 1 Introduction
In this context, wall shear stress (WSS) and oscillatory shear index (OSI) are critical hemodynamic parameters for predicting aneurysm growth and assessing the risk of aneurysm rupture [8]. WSS refers to the tangential force exerted by blood flow on the vessel wall, influencing endothelial cell function and vascular remodeling. Abnormal WSS can contribute to the development and progression of aneurysms by promoting inflammatory responses and weakening the vessel wall structure. OSI, on the other hand, measures the change in the direction of shear stress over the cardiac cycle, indicating areas of disturbed flow that are prone to pathological changes. High OSI values are often associated with regions of the arterial wall that are susceptible to aneurysm formation and growth due to prolonged exposure to oscillatory and low shear stresses. Understanding the spatial and temporal variations of WSS and OSI within aneurysms is essential for identifying high-risk regions and improving the prediction of aneurysm growth and rupture, thereby aiding in developing targeted therapeutic interventions and enhancing patient outcomes.
Traditionally, computational fluid dynamics (CFD) for cardiovascular modeling has relied on methods such as the finite element method (FEM), finite volume method (FVM), and finite difference method (FD). These techniques have been instrumental in providing detailed simulations of fluid flow and structural interactions within the cardiovascular system. However, they often require significant computational resources and can be time-consuming, especially for complex geometries and time-dependent analyses [9, 10, 11].
Recent advancements in computational modeling, particularly the application of Graph Neural Networks (GNNs), offer promising new avenues for cardiovascular blood flow modeling. GNNs provide an innovative approach by leveraging the inherent graph structure of a computational mesh, allowing for efficient and accurate representations of fields defined at mesh vertices, and consequently predictions of hemodynamic parameters.
For instance, Pegolotti et al. [12] utilized MeshGraphNets to predict pressure and flow rates at vessel centerline nodes within a 1D model. Similarly, Suk et al. [13] employed equivariant GNNs on CFD-derived meshes to predict WSS on the surface of single arteries. Their methodology, however, had limitations: the dataset splitting over training and test sets does not reflect any physics- or geometry-based criterion, they did not predict WSS on completely unseen arteries, did not evaluate WSS over time, and did not focus on aneurysms specifically. Furthermore, they primarily used typical machine learning metrics rather than assessing clinically relevant indices such as the OSI. Finally, Suk et al. [14] used GNNs as an efficient black-box surrogate model to estimate 3D velocity fields in the artery lumen.
@@ 2.1.3 Numerical discretization
It is well-known that the flow characteristics in the TAA strongly depend on its shape, which directly influences wall shear indices. Consequently, using patient-specific geometry is essential to obtain accurate results. To achieve this, CT-scan images of a 48-year-old male patient were acquired. Image-processing techniques were then employed to reconstruct a 3D model from the 2D images. Additional branches and extraneous details were removed during this process. Moreover, the outlet boundaries were extended by a specified length in the normal direction to minimize the impact of the outlet boundary condition on the main flow field and enhance numerical stability [22, 23]. The final geometry of the TAA, after these modifications and clean-up, is shown in Fig. 1.
@@
The inflow and outflow boundaries are managed by mass flow inlet and pressure outlet boundary conditions, respectively, based on data provided by [9]. The time-dependent profile for the inlet and both top and bottom outlets is illustrated in Fig. 4.
@@
Several cardiac cycles must pass to achieve fully stationary results. Additionally, to ensure convergence, initial values of zero for pressure and velocity were applied, making the early cycles inaccurate.
@@ 2.5 Graph Neural Networks
Graph Neural Networks (GNNs) [24] [25] [26] are a widely popular class of machine learning algorithms designed to operate on structured data, such as social networks, databases, traffic networks and molecules, suitable to be processed with a graph structure, i.e. as nodes (or vertices) connected by edges that represent the relationship between linked nodes. They have shown outstanding performances in many applications such as traffic flow predictions [27] , molecule property prediction [28] and generation [29], weather forecasting [30].
@@ 3 Results and discussion
Experimental setup
We aim at predicting the WSS and the OSI indices over the surface of a TAA at a specific growth stage by means of GNNs. Our numerical study consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 100%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 75%. This task focus on the interpolation capabilities of the model over shear indices at different growth stages of the TAA.
Training and test loss are evaluated over 5 different seeds, as possible different weight initializations at the beginning of the training have to be taken in account. Each training phase runs for a maximum of 5000 epochs; the learning rate is set as $\lambda=10^{-3}$.
@@
We conducted a preliminar experimental study to determine how to choose the best GNN architecture to be used in the main experimental framework. Among the several GNN modules proposed in literature, we focused on 4 particular modules: GraphConv, Graph Convolutional Networks (GCNs), Graph Isomorphism Networks (GINs) and Graph Transformers. A detailed description of each module can be found in Appendix A.
In order to choose the optimal GNN module, we conduct both a quantitative and a qualitative comparison. For the quantitative comparison, we compute the relative $\infty$-norm error between the TAWSS computed over the reference WSS values and the TAWSS computed over the predicted WSS values (denoted as $\overline{\text{TAWSS}}$):
@@
Numerical results
@@
The WSS prediction on the 75% size TAA by the best GNN with 4 layers over 5 runs is shown in Figure 17 in comparison with the ground truth, at timestamp $t=0.186$, which corresponds to the systolic peak.
The prediction on the normalized OSI at the final timestamp (as defined in Equation 10) by the same architecture is also provided (Figure 18). Additionally, a threshold $\mathsf{thresh}=0.3$ is applied on the normalized OSI, to highlight the accuracy of prediction in high risk areas (Figure 19) [41].
@@
Figure 19: Ground truth (left) vs prediction (right).
The GNN is able to well predict qualitatively the areas where the peaks of the stress indices will be reached in the systolic phase. The good performance of the predictor is additionally highlighted by the node-wise predictions in selected nodes, where the overall trend is well catched modulo the small obscillations.
@@ 4 Conclusion
In this paper we developed a novel framework for the prediction of shear stress indices, such as the WSS, integrating Full Order Models techniques with surrogate models; in particular we obtained high fidelity data from blood flow modeling via Finite Volumes discretization in order to train Graph Neural Networks. This approach overcomes the traditional mesh refinement problem, as GNNs can yield instantaneous evaluations on the input meshes, which can be made of different number of nodes. GNNs exhibit good qualitative performances in predicting WSS and OSI, providing an alternative tool to monitor shear stress peaks at different volume sizes and consequently to quantify the risk of ruptures or dissections in the TAA.
Our method is not free from open issues. Firstly, the individual performances of the nodes need to be enhanced to ensure accurate local predictions. One way to achieve this is by treating the data as temporal sequences and using temporal models, namely, Temporal GNNs [42]. This approach would involve a significant preprocessing stage, where we’d need to define a sliding window of a certain length to create temporal graph sequences that are long enough to capture temporal behavior, yet short enough to be processed by a relatively simple Temporal GNN, in order to avoid issues like vanishing gradients [43]. Another option would be to process the high-resolution data as point clouds, allowing us to apply state-of-the-art methods for prediction tasks on point clouds, such as PointNets [44]. Eventually, with more advanced hardware sources, we could fine-tune our model via grid-search methods to find the optimal configuration.
@@ Acknowledgements
GAD, PCA, and GR acknowledge the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. SS and GR acknowledge the support provided by PRIN ”FaReX - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning” project, and by INdAM-GNCS 2019–2020 projects and PON ”Research and Innovation on Green related issues” FSE REACT-EU 2021 project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them. In addition, the authors would like to acknowledge INdAM–GNCS.

### v2

# Mesh-Informed Reduced Order Models for Aneurysm Rupture Risk Prediction

@@ Abstract
The complexity of the cardiovascular system needs to be accurately reproduced in order to promptly acknowledge health conditions; to this aim, advanced multifidelity and multiphysics numerical models are crucial. On one side, Full Order Models (FOMs) deliver accurate hemodynamic assessments, but their high computational demands hinder their real-time clinical application. In contrast, ROMs provide more efficient yet accurate solutions, essential for personalized healthcare and timely clinical decision-making. In this work, we explore the application of computational fluid dynamics (CFD) in cardiovascular medicine by integrating FOMs with ROMs for predicting the risk of aortic aneurysm growth and rupture. Wall Shear Stress (WSS) and the Oscillatory Shear Index (OSI), sampled at different growth stages of the thoracic aortic aneurysm, are predicted by means of Graph Neural Networks (GNNs). GNNs exploit the natural graph structure of the mesh obtained by the Finite Volume (FV) discretization, taking into account the spatial local information, regardless of the dimension of the input graph. Our experimental validation framework yields promising results, confirming our method as a valid alternative that overcomes the curse of dimensionality.
@@ 1 Introduction
In this context, wall shear stress (WSS) and oscillatory shear index (OSI) are critical hemodynamic parameters for predicting aneurysm growth and assessing the risk of aneurysm rupture [8, 9, 10, 11, 7, 4]. WSS refers to the tangential force exerted by blood flow on the vessel wall, influencing endothelial cell function and vascular remodeling. Abnormal WSS can contribute to the development and progression of aneurysms by promoting inflammatory responses and weakening the vessel wall structure. OSI, on the other hand, measures the change in the direction of shear stress over the cardiac cycle, indicating areas of disturbed flow that are prone to pathological changes. High OSI values are often associated with regions of the arterial wall that are susceptible to aneurysm formation and growth due to prolonged exposure to oscillatory and low shear stresses. Understanding the spatial and temporal variations of WSS and OSI within aneurysms is essential for identifying high-risk regions and improving the prediction of aneurysm growth and rupture, thereby aiding in developing targeted therapeutic interventions and enhancing patient outcomes.
Traditionally, computational fluid dynamics (CFD) for cardiovascular modeling has relied on methods such as the finite element method (FEM), finite volume method (FV), and finite difference method (FD). These techniques have been instrumental in providing detailed simulations of fluid flow and structural interactions within the cardiovascular system. However, they often require significant computational resources and can be time-consuming, especially for complex geometries and time-dependent analyses [12, 13, 14].
Recent advancements in computational modeling, particularly the application of Graph Neural Networks (GNNs), offer promising new avenues for cardiovascular blood flow modeling. GNNs provide an innovative approach by leveraging the inherent graph structure of a computational mesh, allowing for efficient and accurate representations of fields defined at mesh vertices, and consequently, for predictions of hemodynamic parameters.
For instance, Pegolotti et al. [15] utilized MeshGraphNets to predict pressure and flow rates at vessel centerline nodes within a 1D model. Similarly, Suk et al. [16] employed equivariant GNNs on CFD-derived meshes to predict WSS on the surface of single arteries. Their methodology, however, had limitations: the dataset splitting over training and test sets does not reflect any physics- or geometry-based criterion, they did not predict WSS on completely unseen arteries, did not evaluate WSS over time, and did not focus on aneurysms specifically. Furthermore, they primarily focused on typical machine learning metrics to evaluate the quality of the model training, rather than assessing clinically relevant indices such as the OSI. Finally, Suk et al. [17] used GNNs as an efficient black-box surrogate model to estimate 3D velocity fields in the artery lumen.
@@ 2.1.3 Numerical discretization
It is well-known that the flow characteristics in the TAA strongly depend on its shape, which directly influences wall shear indices [25, 26]. Consequently, using patient-specific geometry is essential to obtain accurate results. To achieve this, CT-scan images of a 48-year-old male patient were acquired. Image-processing techniques were then employed to reconstruct a 3D model from the 2D images. The model is adopted from [27], where all details regarding the utilized software and the image processing of the CT data are provided. Additional branches and extraneous details were removed during this process. Moreover, the outlet boundaries were extended by a specified length in the normal direction to minimize the impact of the outlet boundary condition on the main flow field and enhance numerical stability [28, 29]. The final geometry of the TAA, after these modifications and clean-up, is shown in Fig. 1.
@@
The inflow and outflow boundaries are managed by mass flow inlet and pressure outlet boundary conditions, respectively, based on data provided by [12]. The time-dependent profile for the inlet and both top and bottom outlets is illustrated in Fig. 4. Furthermore, only the fluid flow is modeled; thus, there is no vascular structure modeling, and the walls are assumed to be rigid. Additionally, a no-slip condition is applied on the vascular walls.
@@
Three cardiac cycles are required to achieve fully stationary results. Additionally, to ensure convergence, initial values of zero for pressure and velocity were applied, making the early cycles inaccurate.
@@ 2.5 Graph Neural Networks
Graph Neural Networks (GNNs) [30] [31] [32] are a widely popular class of machine learning algorithms designed to operate on structured data, such as social networks, databases, traffic networks, and molecules, suitable to be processed with a graph structure, i.e. as nodes (or vertices) connected by edges that represent the relationship between linked nodes. They have shown outstanding performances in many applications such as traffic flow predictions [33] , molecule property prediction [34] and generation [35], and weather forecasting [36].
@@ 3.1 Experimental setup
We aim at predicting the WSS and the OSI indices over the surface of a TAA at a specific growth stage by means of GNNs. Our numerical study is split in two steps:
  • 1.
    the first step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 100%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 75%. This task focus on the interpolation capabilities of the model over shear indices at different growth stages of the TAA.
  • 2.
    the second step consists of training a GNN over a training set composed by graph snapshots taken from TAAs of sizes 25%,50% and 75%, while it is evaluated over a test set composed by graph snapshots taken from a TAA of size 100%. This task focus on the extrapolation capabilities of the model over shear indices at different growth stages of the TAA. This task is clinically more relevant as we would like to predict the WSS on future (and therefore larger) volume sizes of the TAA.
We choose the Mean Square Error as loss functional, and the Adam algorithm as optimizer, with the learning rate set to $\lambda=10^{-3}$.
@@
We conducted a preliminar experimental study to determine how to choose the best GNN architecture to be used in the main experimental framework. Among the several GNN modules proposed in literature, we focused on 4 particular modules: GraphConv, Graph Convolutional Networks (GCNs), Graph Isomorphism Networks (GINs) and Graph Transformers. A detailed description of each module can be found in A.
In order to choose the optimal GNN module, we conduct both a quantitative and a qualitative comparison on the interpolation task. For the quantitative comparison, we compute the relative $\infty$-norm error between the TAWSS computed over the reference WSS values and the TAWSS computed over the predicted WSS values (denoted as $\overline{\text{TAWSS}}$):
@@
3.2 Ablation study on the number of layers
@@
3.3 Interpolation task: numerical results
The WSS prediction on the 75% size TAA by the best GNN with 4 layers is shown in Figure 17 in comparison with the ground truth, at timestamp $t=0.186$, which corresponds to the systolic peak.
The prediction on the normalized OSI at the final timestamp (as defined in Equation 10) by the same architecture is also provided (Figure 18). Additionally, a threshold $\mathsf{thresh}=0.3$ is applied to the normalized OSI, to highlight the accuracy of the prediction in high-risk areas (Figure 19) [47].
@@ 3.3 Interpolation task: numerical results
Figure 19: OSI at threshold 0.3 over a TAA of 75% size percentage. Ground truth (left) vs prediction (right).
The GNN is able to well predict qualitatively the areas where the peaks of the stress indices will be reached in the systolic phase. The good performance of the predictor is additionally highlighted by the node-wise predictions in selected nodes, where the overall trend is well caught modulo the small oscillations.
3.4 Extrapolation task: numerical results
For the extrapolation task we show the prediction on the normalized WSI (at systolic peak) and OSI (at the final timestamp) with and without thresholds (Figures 20, 21 and 22, respectively).
Similar trends in the error distribution of WSS and OSI are observed in the extrapolation case, similar to those in interpolation. The network effectively predicts the regions of high concentration of WSS (Figure 20) and OSI (Figure 21). The range of relative errors for WSS in both interpolation and extrapolation cases is comparable, with notably higher absolute errors in the aneurysm bleb region. This can be attributed to the complex flow dynamics in that region compared to the aneurysm sac. Specifically, Figure 23 illustrates the velocity vectors at the systolic peak within the aneurysm, highlighting more pronounced vortical structures in the bleb and greater flow uniformity in the sac. This complexity amplifies flow non-linearity, thereby influencing predictions and resulting in higher error values.
While the OSI prediction error distribution is similar for interpolation and extrapolation, a more detailed quantitative analysis reveals that the maximum absolute error in the extrapolation case is 0.08—significantly lower than the interpolation case, which has a maximum error of 0.5, as shown in Figure 18. The error remains concentrated around the aneurysm bleb, again linked to the heightened non-linearity of the flow field in this region. As previously noted, clinical guidelines identify an OSI value exceeding 0.3 as indicative of rupture risk. Based on this criterion, Figure 22 compares the FOM and GNN results, demonstrating a high level of agreement.
Figure 20: Ground truth (left) vs prediction (center) vs relative error percentage (left) on WSS index over a TAA of 100% size percentage, at systolic time. The prediction is made by a GNN with 4 layers. In the figure, the WSS has been rescaled in $[0,0.25]$ to make the comparison clearer.
Figure 21: Ground truth (left) vs prediction (center) vs absolute error (left) on OSI index over a TAA of 100% size percentage. The prediction is made by a GNN with 4 layers.
Figure 22: OSI at threshold 0.3 over a TAA of 100% size percentage. Ground truth (left) vs prediction (right).
Figure 23: Velocity vectors inside aneurysm at the systolic peak.
@@ 4 Conclusion
In this paper we developed a novel framework for the prediction of shear stress indices, such as the WSS, integrating Full Order Models techniques with surrogate models; in particular we obtained high fidelity data from blood flow modeling via Finite Volumes discretization in order to train Graph Neural Networks. This approach overcomes the traditional mesh refinement problem, as GNNs can yield instantaneous evaluations on the input meshes, which can be made of different numbers of nodes. GNNs exhibit good qualitative performance in predicting WSS and OSI, both on interpolation and extrapolation tasks, providing an alternative tool to monitor shear stress peaks at different volume sizes and consequently to quantify the risk of ruptures or dissections in the TAA.
Our method is not free from open issues. Firstly, the implemented GNNs are permutation equivariant, but not SE3 equivariant. Although our results don’t show limitations arising from the lack of this property, the implementation of SE3 equivariance would make our approach more robust to data rotation (see, for instance, the approach used in [48]). The individual performances of the nodes need to be enhanced to ensure accurate local predictions. One way to achieve this is by treating the data as temporal sequences and using temporal models, namely, Temporal GNNs [49]. This approach would involve a significant preprocessing stage, where we’d need to define a sliding window of a certain length to create temporal graph sequences that are long enough to capture temporal behavior, yet short enough to be processed by a relatively simple Temporal GNN, in order to avoid issues like vanishing gradients [50]. Another option would be to process the high-resolution data as point clouds, allowing us to apply state-of-the-art methods for prediction tasks on point clouds, such as PointNets [51]. Eventually, with more advanced hardware sources, we could fine-tune our model via grid-search methods to find the optimal configuration. Another notable limitation of this study is the reliance on data from a single patient. While the results provide valuable insights, they may not fully capture the variability across a broader population. Future work should focus on validating the findings using data from multiple patients to enhance the generalizability of the proposed approach. Expanding the dataset will also allow for a more robust assessment of inter-patient variability and its implications on the model’s predictive performance.
@@ Acknowledgements
G.A.D., P.C.A., and G.R. acknowledge the support provided by the European Union - NextGenerationEU, in the framework of the iNEST - Interconnected Nord-Est Innovation Ecosystem (iNEST ECS00000043 – CUP G93C22000610007) project. S.S. and G.R. acknowledge the support provided by PRIN ”FaReX - Full and Reduced order modelling of coupled systems: focus on non-matching methods and automatic learning” project, and by INdAM-GNCS 2019–2020 projects and PON ”Research and Innovation on Green related issues” FSE REACT-EU 2021 project. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them. In addition, the authors would like to acknowledge INdAM–GNCS.